solve by the system by the elimination method
5x+3y=-11
7x-2y=13
the solution is?

Answer by algebrahouse.com(1659)

solve by the system by the elimination method
5x + 3y = -11-----> 10x + 6y = -22 {multiplied top equation by 2}
7x - 2y = 13------> 21x - 6y = 39 {multiplied bottom equation by 3}

31x = 17 {added the two new equations together}
x = 17/31 {divided both sides by 31}

5x + 3y = -11 {top equation}
5(17/31) + 3y = -11 {substituted 17/31, in for x, into top equation}
85/31 + 3y = -11 {multiplied 5 by 17/31}
3y = -426/31 {subtracted 85/31 from both sides}
y = -142/31

x = 17/31 and y = -142/31
www.algebrahouse.com